# Generating metamers of human scene understanding

### v2

@@ 5.2.3 High-level visual features
DreamSim distance provides the clearest evidence for high-level semantic alignment contributing to a metameric judgment. We found that smaller distance values predict more “same” responses to generations (Figure 5, rightmost plot). This is likely because DreamSim was specifically trained on human judgments using a two-alternative forced-choice paradigm to capture human-like notions of visual similarity. CLIP also predicted this relationship between semantic alignment and the likelihood of a generation becoming a metamer (Figure 5). However, metamerism rate increased with semantic alignment only when generations were based on the viewer’s own fixations. When generations were from random fixations, higher CLIP similarity did not predict more “same” metameric judgments. We suggest that this interaction may be due to random fixations often falling on contextually irrelevant regions, thereby exposing semantic details that are not aligned with the viewer’s internal scene representation. A similar interaction between own-fixation and random conditions appears in DreamSim, but only at high distances. Together, we interpret our results to mean that scenes generated from a viewer’s attention have better high-level semantic alignment with the viewer’s internal scene representation and are therefore more likely to be confused as originals. Additional object-level visual feature analyses can be found in Appendix A.4.
@@ 5.3 Foveal and peripheral features both contribute to metameric judgments
We ran an ablation experiment to isolate the contribution of foveal and peripheral conditioning in MetamerGen. We recruited 10 additional viewers to participate in a same-different task similar to the primary experiment reported above, but with four second-image conditions that systematically assessed the impact of conditioning: identical original images (true “same” responses), generated images using both foveal and peripheral conditioning (as in the primary experiment), generated images using peripheral-only, and generated images using foveal-only conditioning. We found that while both foveal and peripheral conditioning played a role in whether a generation becomes a metamer, the role played by peripheral conditioning was greater. As expected, the full model had the highest metamer rate of 54.5%, compared to the second highest rate of 45.8% in the peripheral-only generation condition. The metamer rate in the foveal-only condition was only 8.4%. This much lower rate was due to the generations in that condition correctly capturing fixated details but differing from originals substantially in the non-fixated regions. In contrast, generations in the peripheral-only condition were more likely to be judged metamers because they better captured the global scene structure and layout. Nevertheless, this ablation analysis shows that the additional conditioning from foveal inputs, when combined with the peripheral conditioning, contributes visual and semantic information that leads togenerations that are better aligned with human scene understanding than generations produced from peripheral conditioning alone.
We replicated the multi-level visual feature analysis from Section 5.2 for each condition of this ablation experiment and found that: (1) under equivalent levels of feature similarity, generations from the full model are most likely to produce metamers, and (2) in general, feature similarity predicted participant judgments the best among full model generations, more poorly among peripheral-only generations, and very little among foveal-only generations. For more detailed results and figures, see Appendix  A.9.
@@ 7 Discussion
This paper introduces MetamerGen, a latent diffusion model for generating images that are aligned with human scene representations enough to become scene metamers. MetamerGen demonstrates that sparse conditioning based on peripheral gist and fixation-specific details is sufficient to produce human-aligned scene representations. The work described in this paper should be of broad interest to communities ranging from cognitive scientists to machine learning researchers. For cognitive scientists, MetamerGen becomes a powerful tool for exploring the human representation of scenes. It enables the testing of fixation-specific hypotheses about what a person’s scene understanding would be given their allocation of attention during scene viewing. For machine learning researchers, MetamerGen advances generative modeling by producing from sparsely sampled inputs semantically coherent scenes that are demonstrated to align with human scene representations. This greater human alignment will contribute to next-generation assistive technologies capable of interacting more naturally with humans.
@@
This behavioral experiment presented in this work was conducted in accordance with ethical guidelines for a human subjects research. The study protocol was reviewed and approved by the Institutional Review Board.
All participants provided informed consent before participating in the behavioral experiment. Participants were fully informed about study procedures during trials. Participation was entirely voluntary, with participants retaining the right to withdraw at any time without penalty. The data we collected involved non-invasive eye-tracking using the EyeLink 1000 eye-tracker. All data were de-identified (codified) and cannot be linked back to individual participants.
@@ Acknowledgments
We would like to thank the National Science Foundation for supporting this work through awards 2123920 and 2444540 to GJZ, and the National Institutes of Health through their award R01EY030669, also to GJZ. AL is supported by NSF-GRFP award 2234683. AG is supported by NSF grants IIS-2123920, IIS-2212046 awarded to DS. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agency.
@@ Appendix A Appendix
A.1 Eye-tracking behavioral paradigm
A.1.1 Eye-tracking methodology and fixation coordinate extraction
Eye movements were recorded using an EyeLink 1000 eye-tracker (SR Research Ltd., 2006) configured with the Tower Mount setup. This configuration positions the infrared camera above the participant via a mirror, providing an unobstructed view while enabling monocular tracking across $55^{\circ}$ horizontally and $45^{\circ}$ vertically. Participants viewed stimuli on a 27 inch $2560\times 1440$ resolution 240Hz OLED monitor positioned $24$ inches from their eyes (subtending approximately $55^{\circ}\times 30^{\circ}$ visual angle). Prior to each experimental session, a standard $13$-point calibration procedure was performed to ensure accurate gaze tracking. During free-viewing trials, fixations were detected online using the EyeLink’s built-in saccade detection algorithm.
A.1.2 Experimental design overview
Figure 6: Real-time metameric judgment paradigm. Each trial begins with drift correction and central fixation, followed by free viewing of an original scene for a predetermined number of fixations. After image offset, participants maintain central fixation for 5 seconds while fixation coordinates are transmitted via API to MetamerGen for a real-time image generation. The generated image (or original as control) is then presented for 200ms, followed by a same-different judgment using a gamepad within a 10-second response window.
During a given trial, given fixation coordinates $(x,y)$ from eye-tracking data, we map each fixation to the corresponding patch token in DINOv2’s $32\times 32$ grid. For $448\times 448$ input images, each patch token represents a $14\times 14$ pixel region (roughly $1.2^{\circ}\times 1.2^{\circ}$ visual angle). Fixation coordinates are normalized to this grid space, with the nearest patch token selected and all others zeroed out, forcing the model to reconstruct the entire scene from sparse fixation inputs.
@@ A.3 Perceiver-based Resampler Architecture
A.4 Additional visual feature drivers of metameric judgment
Figure 7: Object detection errors predict metameric perception: (Top) mAP scores demonstrate that higher precision accuracies (from mAP 60% to mAP 80%) with better alignment at strict localization boundaries correlate with increased ”same” metameric judgments. (Bottom) Object detection metrics show a positive relationship where improvements in model precision, recall, and F1 scores correspond to increased ”same” metameric judgments.
Multi-object recognition & localization
@@ A.6 Additional generation visualizations based on fixated inputs
Figure 9: Additional metameric vs. non-metameric judgment example images based on human fixations. (Left) Original images with human fixations overlaid in red and corresponding generated images judged as ”same” by participants. (Right) Original images with fixations and generated images judged as ”different” by participants.
Figure 10: Additional metameric vs. non-metameric judgment example images based on randomly-sampled fixations. (Left) Original images with randomly-sampled fixations overlaid in red and corresponding generated images judged as ”same” by participants. (Right) Original images with fixations and generated images judged as ”different” by participants.
@@ A.9 Foveal and peripheral token contributions to metameric judgments
We present further statistics and analyses of the behavioral ablation experiment presented in Section 5.3.
Figure 13 provides qualitative examples of generations from each treatment that were judged “same” and “different”, together with a plot of the fool rate in each condition. The generations in the full-model condition are both qualitatively the highest-quality and quantitatively the most likely to fool participants.
@@ A.10 DINOv2 versus CLIP as the vision encoder of MetamerGen
We observed that CLIP features perform significantly worse at encoding peripheral information, reflected by the higher FID values obtained in Figure 15 (left),  16. In contrast to DINOv2 tokens, the FID value does not show a significant decrease when we reduce the blur level, meaning that CLIP encodes similar (impoverished) information in its patch tokens for the blurred and non-blurred images, focusing on global scene category, at the expense of more fine-grained scene structure. This is consistent with previous observations made regarding CLIP’s relatively low ability to encode contextual information with its patch tokens (Li et al., 2025). For foveal information, increasing the number of tokens has a similar effect for both encoders, though a single DINOv2 token seems to encode more information than a single CLIP token. Overall, our choice of DINOv2 is mainly motivated by its ability to accurately encode both the peripheral and foveal information present in the image.

### v3

@@ 5.2.3 High-level visual features
DreamSim distance provides the clearest evidence for high-level semantic alignment contributing to a metameric judgment. We found that smaller distance values predict more “same” responses to generations (Figure 5, rightmost plot). This is likely because DreamSim was specifically trained on human judgments using a two-alternative forced-choice paradigm to capture human-like notions of visual similarity. CLIP also predicted this relationship between semantic alignment and the likelihood of a generation becoming a metamer (Figure 5). However, metamerism rate increased with semantic alignment only when generations were based on the viewer’s own fixations. When generations were from random fixations, higher CLIP similarity did not predict more “same” metameric judgments. We suggest that this interaction may be due to random fixations often falling on contextually irrelevant regions, thereby exposing semantic details that are not aligned with the viewer’s internal scene representation. A similar interaction between own-fixation and random conditions appears in DreamSim, but only at high distances. Together, we interpret our results to mean that scenes generated from a viewer’s attention have better high-level semantic alignment with the viewer’s internal scene representation. Additional object-level visual feature analyses are discussed in Appendix A.5.
@@ 5.3 Foveal and peripheral features both contribute to metameric judgments
We ran an ablation experiment to isolate the contribution of foveal and peripheral conditioning in MetamerGen. We recruited 10 additional viewers to participate in a same-different task similar to the primary experiment reported above, but with four second-image conditions that systematically assessed the impact of conditioning: identical original images (true “same” responses), generated images using both foveal and peripheral conditioning (as in the primary experiment), generated images using peripheral-only, and generated images using foveal-only conditioning. We found that while both foveal and peripheral conditioning played a role in whether a generation becomes a metamer, the role played by peripheral conditioning was greater. As expected, the full model had the highest metamer rate of 54.5%, compared to the second highest rate of 45.8% in the peripheral-only generation condition. The metamer rate in the foveal-only condition was only 8.4%. This much lower rate was due to the generations in that condition correctly capturing fixated details but differing from originals substantially in the non-fixated regions. In contrast, generations in the peripheral-only condition were more likely to be judged metamers because they better captured the global scene structure and layout. Nevertheless, this ablation analysis shows that the additional conditioning from foveal inputs, when combined with the peripheral conditioning, contributes visual and semantic information that leads to generations better aligned with human scene understanding than generations produced from peripheral conditioning alone. We replicated the multi-level visual feature analysis from Section 5.2 for each condition of this ablation experiment in Appendix  A.9.
@@ 7 Discussion
This paper introduces MetamerGen, a latent diffusion model for generating images that are aligned with human scene representations enough to become scene metamers. MetamerGen demonstrates that sparse conditioning based on peripheral gist and fixation-specific details is sufficient to produce human-aligned scene representations, and should be of broad interest to communities ranging from cognitive scientists to machine learning researchers. For cognitive scientists, MetamerGen becomes a powerful tool for exploring the human representation of scenes. It enables the testing of fixation-specific hypotheses about what a person’s scene understanding would be given their allocation of attention during scene viewing. For machine learning researchers, MetamerGen advances generative modeling by leveraging sparsely sampled inputs to produce semantically coherent scene generations that are broadly aligned with human scene representations. This greater human alignment will contribute to next-generation assistive technologies capable of interacting more naturally with humans. Lastly, while we found that attention-guided conditioning yielded stronger correlations across all feature hierarchies (Figs. 4, 5), we also found no significant difference in metamerism rates between random and attention-guided conditions and this has practical implications for the collection of large-scale datasets of scene metamers because it removes the necessity for eye-data collection.
@@
The behavioral data presented in this work was collected in accordance with ethical guidelines determined by the review board at Stony Brook University responsible for overseeing research on human subjects. All subjects provided informed consent before participating in the behavioral experiment in accordance with board-approved protocols and forms. Participation was entirely voluntary, and participants were free to withdraw at any time without penalty. All behavioral data collected for this study, including the gaze-fixation data collected using the EyeLink 1000 eye-tracker, were de-identified (codified) prior to analysis so that there was no link back to the individual participants.
@@ Acknowledgments
We would like to thank the National Science Foundation for supporting this work through awards 2123920 and 2444540 to GJZ, and the National Institutes of Health through their award R01EY030669, also to GJZ. AL is supported by NSF-GRFP award 2234683. AG is supported by NSF grants IIS-2123920, IIS-2212046 awarded to DS. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agency.
@@ Appendix A Appendix
A.1 Experiment details
Figure 6: Real-time paradigm for determining scene metamers. Each trial begins with drift correction and central fixation, followed by free viewing of an original scene for a predetermined number of fixations. After image offset, participants maintain central fixation for 5 seconds while fixation coordinates are transmitted via API to MetamerGen for real-time image generation. The generated image (or original, depending on the condition) is then presented to the viewer for 200ms, followed by an enforced same-different behavioral judgment via a gamepad within a 10-second response window.
For each behavioral trial, we map the fixation coordinate $(x,y)$ of each viewing fixation to the corresponding patch token in DINOv2’s $32\times 32$ grid. For $448\times 448$ input images, each patch token represents a $14\times 14$ pixel region (roughly $1.2^{\circ}\times 1.2^{\circ}$ visual angle). Fixation coordinates are normalized to this grid space, with the nearest patch token selected and all others zeroed out, forcing the model to reconstruct the entire scene from sparse fixation inputs.
Eye movements were recorded using an EyeLink 1000 eye-tracker (SR Research Ltd., 2006) in tower-mount configuration. This configuration positions the infrared camera above the participant via a mirror, providing an unobstructed view while enabling monocular tracking over a wide visual angle ($55^{\circ}$ horizontal and $45^{\circ}$ vertical). The participant, whose head was fixed by a chin rest and head restraint, viewed stimuli on a 27 inch ($2560\times 1440$ resolution) 240Hz OLED monitor positioned $24$ inches from their eyes, which subtended approximately $55^{\circ}\times 30^{\circ}$ degrees of visual angle on their retina. Prior to each experimental session, a standard $13$-point calibration procedure was performed to ensure accurate eye-gaze tracking. Default algorithms were used to detect eye movements and fixations during both the fixation-dependent free viewing of scenes and all subsequent analyses of eye-gaze fixations.
@@ A.3 Perceiver-based Resampler Architecture
A.4 Stepwise regression model details
A.5 Additional drivers of metameric judgment
A.5.1 Low-level texture features
As discussed in Section 5.2, we found that greater oriented edge contrast in the generated image (as measured by Gabor filters) predicts more “same” judgments. Figure 7 presents the results of this analysis visually, demonstrating the effect for 0°, 45°, and 135° filters. (The effect was not significant for the 90°horizontal filter.) The right side of the figure visualizes specific generated images with high or low filter responses at each orientation.
A.5.2 Multi-object recognition & localization
Figure 8: Object detection errors predict metameric perception: (Top) mAP scores demonstrate that higher precision accuracies (from mAP 60% to mAP 80%) with better alignment at strict localization boundaries correlate with increased “same” metameric judgments. (Bottom) Object detection metrics show a positive relationship where improvements in model precision, recall, and F1 scores correspond to increased “same” metameric judgments.
@@ A.6 Additional generation visualizations based on fixated inputs
Figure 9: Additional metameric vs. non-metameric judgment example images based on human fixations. (Left) Original images with human fixations overlaid in red and corresponding generated images judged as “same” by participants. (Right) Original images with fixations and generated images judged as “different” by participants.
Figure 10: Additional metameric vs. non-metameric judgment example images based on randomly-sampled fixations. (Left) Original images with randomly-sampled fixations overlaid in red and corresponding generated images judged as “same” by participants. (Right) Original images with fixations and generated images judged as “different” by participants.
@@ A.9 Foveal and peripheral token contributions to metameric judgments
We present further statistics and analyses of the behavioral ablation experiment presented in Section 5.3. Figure 13 provides qualitative examples of generations from each treatment that were judged “same” and “different”, together with a plot of the fool rate in each condition. The generations in the full-model condition are both qualitatively the highest-quality and quantitatively the most likely to fool participants.
@@ A.10 DINOv2 versus CLIP as the vision encoder of MetamerGen
We observed that CLIP features perform significantly worse at encoding peripheral information, reflected by the higher FID values obtained in Figure 15 (left), and Figure 16. In contrast to DINOv2 tokens, the FID value does not show a significant decrease when we reduce the blur level, meaning that CLIP encodes similar (impoverished) information in its patch tokens for the blurred and non-blurred images, focusing on global scene category, at the expense of more fine-grained scene structure. This is consistent with previous observations made regarding CLIP’s relatively low ability to encode contextual information with its patch tokens (Li et al., 2025). For foveal information, increasing the number of tokens has a similar effect for both encoders, though a single DINOv2 token seems to encode more information than a single CLIP token. Overall, our choice of DINOv2 is mainly motivated by its ability to accurately encode both the peripheral and foveal information present in the image.